ABC and VDC are two isosceles triangles drawn on both sides of BC .E is the symmetric of C with respect to A and F is the symmetric of C with respect to D.
Show that E,B and F are collinear
1
Expert's answer
2020-10-27T17:25:15-0400


Since "E" is symmetric With respect to "A" , "AC=AE" , similarly to "CB=AB" .

Let's lower the medians from the vertices "A" and "D" to the base "BC" , these will be the medians "AH" and "DH" , so "AD" is perpendicular to "CB" and divides it in half. "AD" is the middle line of "\triangle ECF" (since "A" and "D" are the midpoints of "EC" and "CF" , respectively), hence "CB" is perpendicular to "EF", since the ratio of the triangles "\frac{\triangle CAD}{\triangle CEF} = \frac{1}{2}" , and "\frac{CH}{CB}=\frac{1}{2}" , then B belongs to the straight "EF" , hence "B, E, F" are collinear.